Solve by the Elimination Method.
7r - 5s = 11.
5r + 7s = 103.

Answer by checkley77(12844)

7r - 5s = 11. multipl by 7
5r + 7s = 103. multiply by 5 & add.
49r-35s=77
25r+35s=515
--------------
74r=592
r=592/74
r=8 ans.
7*8-5s=11
56-5s=11
-5s=11-56
-5s=-45
s=-45/-5
s=9 ans.
Proof:
5*8+7*9=103
40+63=103
103=103